OpenStudy (anonymous):

help plz Solve the equation. x – 2/7 = 3 A. x = –3 2/7 B. x = –2 5/7 C. x = 2 5/7 D. x = 3 2/7

OpenStudy (hhmorris98):

D

OpenStudy (mathstudent55):

You want x alone. 2/7 is being subtracted from x. You need to add 2/7 to both sides. x - 2/7 + 2/7 = 3 + 2/7

OpenStudy (mathstudent55):

The -2/7 and + 2/7 on the left side add to 0 leaving just x on the left side. On the right side, you add 3 and 2.7 getting 3 2/7 The solution is x = 3 2/7

OpenStudy (anonymous):

Solve. Express the answer in simplest form. https://static.k12.com/bank_packages/files/media/mathml_d1fea9ed23abaef0c4d6d34c985e6440b88dbb15_1.gif w =

OpenStudy (anonymous):

@texaschic101

OpenStudy (anonymous):

@MathiznotEZ

OpenStudy (mathstudent55):

In this case, 1/5 is being added to w. You want w alone, so you need to subtract 1/5 Since you must do the same to both sides of an equation, you need to subtract 1/5 from both sides.

OpenStudy (mathstudent55):